OpenStudy (anonymous):

What is the simplified form of (10p^2 over m^2 + m – 6) over (2p over 3m – 6)

OpenStudy (espex):

Is this your problem?\[\frac{\frac{10p^{2}}{m^{2}+m-6}}{\frac{2p}{3m-6}}\]

OpenStudy (anonymous):

yes

OpenStudy (espex):

Recall that dividing two fractions is the same as multiplying the numerator by the inverse of the denominator.

OpenStudy (espex):

So flip the denominator and multiply. \[(\frac{10p^{2}}{m^{2}+m-6})({\frac{3m-6}{2p}})\]